log5 (base of 5) of x + 3 (no parentheses) = log5 (base of 5) of (x-20) + 4
Answer by stanbon(75887)

log5 x + 3 = log5 (x-20) + 4
log5 x - log5 (x-20) = 1
log5 [x/(x-20)] 1
x/(x-20) = 5
x = 5(x-20)
x = 5x - 100
4x=100
x = 25
------------
Cheers,
Stan H.